OpenStudy (cedric):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set1: (4,4) and (2, 6) Set 2: (2,4) and (4,2) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (cedric):

i thiink its B

OpenStudy (anonymous):

To find slope, you use (y-y)/(x-x) So, i'll do set one: (4-6)/(4-2) SO slope=1 DO the same for set two and compare!:) Hope I helped!:D

OpenStudy (cedric):

So the second one would be set up like (2-2) / (4-4) slope = 0?

OpenStudy (cedric):

no no no it be 1 also because 4-2= 2 and 2/2 equals 1

OpenStudy (anonymous):

No sir. It would be (4-2)/(2-4)..

OpenStudy (cedric):

so now (4-2) = 2 and (2-4) = -2 divide those and get -1

OpenStudy (anonymous):

Yessir:) So, which statement is true:)

OpenStudy (cedric):

B :D